OpenStudy (anonymous):

Corinne travels from home to work at an average speed from 50 miles per hour and returned home by the same route at 60 miles per hour it takes her more than 10 minutes to get to work then it takes her to get home how many miles is it from corinnes home to work

OpenStudy (anonymous):

a 25 b 35 c 50 d 75 e 90

OpenStudy (perl):

you can use distance = rate * time home to work : d1 = 50 * (t +10/60) work to home : d2 = 60 * t We are given that d1 = d2 , same distance traveled. substitute 50(t+10/60) = 60t solve 50t + 500/60 = 60t once you solve for time, plug back in to find

OpenStudy (perl):

note that the units of speed is miles per hour, so i had to change the 10 minutes to 10/60 hour

OpenStudy (anonymous):

thank you